David recently graduated with a B.A. in economics and was offered a job with a large company that pays $44,000 per year. At about the same time, David’s grandfather died and left him an inheritance of $45,000. David decided to pass up the job, and using $39,000 from the inheritance, he purchased a tool rental business. (David put the other $6,000 into a savings account that pays 9 percent annual interest.) He put his full effort into the new business, and in the first three months of operation he had total sales revenue of $29,500 and total explicit costs of $14,260.

a) What is David’s accounting profit for the period?

$


b) What is David’s economic profit for the period?

$
Answer #1

Answer:

a) What is David’s accounting profit for the period?

Calculation of accounting profit=total revenue-explicit cost

Total revenue=sales revenue + interest from $6000=29500+$540=$30,040

Explicit cost=$14260

Accounting profit=$30040-$14260=$15780


b) What is David’s economic profit for the period?

Economic profit=total revenue-explicit cost-implicit cost

Total revenue=sales revenue + interest from $6000=29500+$540=$30,040

Explicit cost=$14260

Implicit cost=income might be earned if joined a company=44000/4=11000

Economic profit=$30040-14260-$11000=$4780